OpenStudy (anonymous):

An isosceles trapezoid has constant bases of 6 and 12 inches, respectively. Find the approximate change in its area when the equal sides change from 5 to 5.2 inches.

OpenStudy (mayankdevnani):

previously height was sqrt(5^2-3^2)=4 so area=.5*(6+12)*4=36 sq cm now height=sqrt(5.2^2-3^2)=4.25; so area=.5*(12+6)*4.25=38.25

OpenStudy (mayankdevnani):

@rainenzo15